OpenStudy (anonymous):

Can someone please explain to me how to factor this s^2+13s+40?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

will u give fan and medal

OpenStudy (anonymous):

yes if you help me

OpenStudy (mathstudent55):

You have a trinomial of the form: \(x^2 + ax + b\) The variable in your case is s instead of x, but the form is the same. The important thing to see is that the square variable is only multiplied by 1.

OpenStudy (anonymous):

8+5=13 8*5=40

OpenStudy (mathstudent55):

All you need to do is set up two sets of parentheses. \((s~~~~~)(s~~~~~~~)\) Then you need to find two numbers whose product is b and whose sum is a.

OpenStudy (mathstudent55):

If you compare your polynomial with \(x^2 + ax + b\) \(s^2 + 13s + 40\) you have a = 13 and b = 40 What are two numbers that multiply to 40 and add to 13?

OpenStudy (mathstudent55):

40 = 40 * 1 = 40; 40 + 1 = 41 (not 13) 20 * 2 = 40; 20 + 2 = 22 (not 13) 10 * 4 = 40; 10 + 4 = 14 (not 13) 8 * 5 = 40; 8 + 5 = 13 (just what we need) That means the two numbers you need that multiply to 40 and add to 13 are 8 and 5. You already had the two sets of parentheses with s in them, now just write each of the numbers in a parentheses: \(s^2 + 13s + 40 = \) \(=(s~~~~~)(s~~~~~~) \) \(=(s + 8)(s + 5) \)

OpenStudy (mathstudent55):

Keep in mind that this method only works for a square trinomial of the form \(x^2 + ax + b\) where the x^2 term has no coefficient (other than 1).

OpenStudy (mathstudent55):

@riley101 Do you understand?

OpenStudy (mathstudent55):

Would you like to see another example?

OpenStudy (anonymous):

I think i understand but would it be different if it was subtraction instead of addition?

OpenStudy (anonymous):

so like 2s^2-14s-156?

OpenStudy (mathstudent55):

The method is still the same. Let me do this simpler example. Then I'll show you how to do yours. Here is an example where there is a subtraction. Factor: \(x^2 -2x - 15\)

OpenStudy (mathstudent55):

Here you have again a trinomial of second degree. You see an x^2 term with no number multiplying it. It is followed by an x-term and by a number. It follows the form \(x^2 + ax + b\), so we can use our method. 1. Setup two sets of parentheses with just x on the left side \(x^2 - 2x - 15\) \(= (x~~~~~~~)(x~~~~~~~)\)

OpenStudy (mathstudent55):

Now you need two numbers that multiply to -15 and add to -2.

OpenStudy (mathstudent55):

Since -15 is a negative number, the only two numbers that multiply to a negative number are a positive number and a negative number. What two numbers multiply to -15 and add to -2?

OpenStudy (anonymous):

whats the easiest way to find those numbers without having to divide starting from one going up?

OpenStudy (anonymous):

5-3?

OpenStudy (anonymous):

or -3-5?

OpenStudy (mathstudent55):

5 and -3 do multiply to -15. Now add them. What do you get?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

but its not negative

OpenStudy (mathstudent55):

Yes. We need the sum to be -2, not positive 2. All you need is a small adjustment. Instead of 5 and -3, you need -5 and 3.

OpenStudy (anonymous):

okay so (x+-5)(x+3)

OpenStudy (mathstudent55):

Now look: -5 * 3 = -15; -5 + 3 = -2 Now we get the product of -15 and the sum of -2. That confirms that the numbers we need are -5 and 3. Now you just place each number inside a set of parentheses following the x. \((x - 5)(x + 3)\)

OpenStudy (mathstudent55):

You are correct. It's just that x + (-5) is simply x - 5.

OpenStudy (mathstudent55):

Now we can try your problem. Are you ready?

OpenStudy (anonymous):

Okay thank you. I'm ready

OpenStudy (mathstudent55):

Factor \(2s^2-14s-156\)

OpenStudy (mathstudent55):

The first rule of factoring: Try to factor a common factor from all terms if possible.

OpenStudy (mathstudent55):

Do the numbers 2, -14, and 156 have a common factor?

OpenStudy (anonymous):

the common factor is 2

OpenStudy (mathstudent55):

Correct. We first factor a 2 from all terms. Can you do that?

OpenStudy (anonymous):

-7, 78

OpenStudy (anonymous):

is it -156, actually?

OpenStudy (mathstudent55):

When we factor out a common factor, we are using the distributive property in reverse. The result of factoring out a common term must be a factored expression that when you use the distributive property on it, you get back the original expression. \(2s^2-14s-156\) \(2(s^2 - 7s - 78)\)

OpenStudy (mathstudent55):

Notice that if you were to start with the second line above and distribute the 2, you'd end up with your original problem.

OpenStudy (anonymous):

Okay thats what I had initially and i thought that was the final answer lol

OpenStudy (mathstudent55):

Now that we have factored out the 2, we end up with a second degree trinomial that starts with simply \(s^2\). That is the kind of trinomial we have been factoring in this post, so we can do it.

OpenStudy (mathstudent55):

Now we ask the question: What two numbers multiply to -78 and add to -7?

OpenStudy (anonymous):

The thing i am having the hardest time with is finding numbers that are the sum and product of the negative numbers

OpenStudy (mathstudent55):

Start by finding any numbers that multiply to 78. Just think of any two positive numbers that multiply to 78. One pair is always easy: 78 and 1. Can you think of other pairs of numbers whose product is 78?

OpenStudy (mathstudent55):

What is 78 divisible by?

OpenStudy (anonymous):

6 x13

OpenStudy (anonymous):

-13 +6

OpenStudy (mathstudent55):

Excellent. 6 * 13 = 78. Since we really need the product to be -78, and not 78, we need to have a positive number and a negative number. You can have -6 * 13 or 6 * (-13) Both have a product of -78 that we need. Does either one of those two pairs add to -7?

OpenStudy (mathstudent55):

Oh, you got it already. You're too fast for me.

OpenStudy (mathstudent55):

Great. The numbers are -13 and 6 since -13 * 6 = -78 -13 + 6 = -7

OpenStudy (anonymous):

so its 2(s-13)(s+6)?

OpenStudy (mathstudent55):

Now all you need to do is place the two numbers you found int he parentheses like we did above. \(2s^2-14s-156\) \(2(s^2 - 7s - 78)\) \(2(s - 13)(s + 6) \)

OpenStudy (mathstudent55):

Correct. Once again, you beat me to it. Next time, I'll ask you for help!

OpenStudy (anonymous):

THANK YOU SO MUCH!!!

OpenStudy (mathstudent55):

Ok, sorry, but gtg.

OpenStudy (mathstudent55):

BTW, this method works for second degree trinomials in which the coefficient of the square term is 1. If the coefficient of the square term is not 1, there is a different method that is a little more complicated, but it's still pretty simple to do. For example: \(6x^2 - 11x - 35\) cannot be factored using this method. 6, -11, and -35 have no common factors, so you can't factor out a common factor to make this trinomial start with a coefficient of 1.